OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (wolf1728):

Let's use body temperature for a comparison 98.6 F = 37 C Fahrenheit = (9/5)*Centigrade +32 F = 9/5 * 37 + 32 F = 66.6 + 32 F = 98.6 Centigrade = (Fahrenheit -32) * (5/9) C = (98.6 -32) * (5/9) C = 66.6 * 5/9 C = 37

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

Showed my age there - Centigrade is an old term for Celsius. :-)